OpenStudy (anonymous):

find the zeros of this function y= f(x) (y=zero) y=(x+2)(x-1) How do I do this!?

OpenStudy (radar):

You can do this. If a * b = 0 What do we know about a or b or maybe even both of them?

OpenStudy (anonymous):

The zero's are simply -2 and 1

OpenStudy (radar):

What you need to know is that if the product of two factors is zero, then either one or both of those factors must be a zero. then either (x+2) = 0 or (x-1)=0, or both then as LagrangeSon has so elequently stated: x+2=0, then x=-2 and (x-1)=0, then x=1.